# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v2

@@ 5 Experimental Results and Analysis
To verify whether the generated code is not only structurally sound but also executable with minimal intervention, we conduct a manual debugging analysis on five representative papers. For each case, we attempts to execute the generated repository and records the number of lines modified to achieve a successful run. Table 7 describes the total number of lines, the modified lines required, and the percentage of changes per repository. Remarkably, we find that on average, only 0.48% of the code needs to be altered for successful execution. Most modifications involve routine fixes such as updating deprecated OpenAI API calls to their latest versions or correcting simple type conversions. These findings indicate that our system does not merely produce syntactically correct or superficially plausible code, but outputs repositories that are functionally robust and executable with minimal human correction.
6 Conclusion and Limitations
6.1 Conclusion
In this work, we introduce PaperCoder, the framework designed to automatically generate code repositories from machine learning research papers. The framework follows a three-stage pipeline planning, analysis, and coding. The planning phase includes the construction of an overall plan, architectural design via diagrams, core logic planning with file dependencies, and configuration file generation. This plan is then refined through a detailed analysis of per-file implementation requirements, followed by sequential code generation guided by the planning outputs.
We evaluate PaperCoder on a comprehensive experimental Paper2Code benchmark consisting of 90 recent papers from top-tier machine learning conferences, as well as the PaperBench Code-Dev benchmark. Our framework consistently outperforms existing baselines across model-based evaluation metrics. In addition, we conduct human evaluations with graduate-level computer science researchers, who assess the quality of repositories generated for their own papers. Results show that 77% of participants preferred PaperCoder’s implementation over alternatives, and 83% found the outputs practically useful for real-world usage. These findings collectively demonstrate the effectiveness of PaperCoder in bridging the gap between research ideas and working code.
6.2 Limitations
While our PaperCoder shows strong performance on recent machine learning papers, its current scope is limited to this specific domain. Extending the framework to support a broader range of scientific fields remains an important direction for future work. Additionally, our primary evaluation relies on model-based metrics, which may not fully capture executability or runtime correctness. Although we perform case studies involving manual debugging to validate executability, a scalable and automated approach to execution-based evaluation—including fault localization and debugging—is a promising avenue for future research. Incorporating such capabilities would further enhance PaperCoder’s utility in real-world research workflows.
@@ A.1 Implementation Details
All experiments are conducted using the o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we utilize the openreview_scraper [^2] https://github.com/pranftw/openreview_scraper and the Semantic Scholar API [^3] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library [^4] https://github.com/allenai/s2orc-doc2json.
@@ A.2 Human Evaluation Process
To conduct evaluations with human judges, we recruited 13 researchers from the South Korea, majoring in computer science, each with a minimum of 1 published papers. For annotation, they were provided with a 4-page document, which includes the task instructions, annotation examples, and 10 generated repositories grouped into three sets. Each repository was anonymized using a repoX naming format to prevent bias regarding the generation method. Following the question guidelines in the document, annotators reviewed and evaluated the repositories generated by different methods and models. Also, on average, evaluating 10 repositories for a single paper took approximately 45 minutes. Table 13 is a detailed annotation example.
@@ A.3 Reference-Based Evaluation
In the reference-based evaluation setting, the repository may exceed the model’s context length. Following Starace et al. [40], when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
@@ A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the benchmark used for evaluation requires a reproduce.sh script. To meet this requirement, we additionally prompt the coding agent to generate a reproduce.sh file specifically for this benchmark, and we evaluate the repository including this file.
@@
Modified lines 2 1 1 25 10 6.5
@@
Percentage 0.18 0.05 0.06 1.87 0.73 0.48
Table 8: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators selected the repositories generated by our framework as their top choice.
@@
8 6 5 4 2 3
@@
Prompt for analyzing [System] You are an expert researcher, strategic analyzer and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to conduct a comprehensive logic analysis to accurately reproduce the experiments and methodologies described in the research paper. This analysis must align precisely with the paper’s methodology, experimental setup, and evaluation criteria. 1. Align with the Paper: Your analysis must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present your analysis in a logical, well-organized, and actionable format that is easy to follow and implement. 3. Prioritize Efficiency: Optimize the analysis for clarity and practical implementation while ensuring fidelity to the original experiments. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. REFER TO CONFIGURATION: Always reference settings from the config.yaml file. Do not invent or assume any values—only use configurations explicitly provided. [User]
@@
Figure 9: Prompts for analyzing. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design.
@@
Table 9: List of ICLR 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
@@ A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Table 10: List of NeurIPS 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
@@
Table 11: List of ICML 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
Oral 4
Autoformalizing Euclidean Geometry
Poster 4
Recurrent Distance Filtering for Graph Representation Learning
Poster 4
Poster 3.9
Token-level Direct Preference Optimization
Poster 3.9
Oral 3.8
CurBench: Curriculum Learning Benchmark
Poster 3.8
Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment
Poster 3.7
Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments
Poster 3.7
Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge
Poster 3.7
Poster 3.7
Poster 3.7
Poster 3.7
Poster 3.7
Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model
Oral 3.6
Oral 3.6
Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting
Oral 3.6
Poster 3.6
Poster 3.6
Graph Distillation with Eigenbasis Matching
Poster 3.6
Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?
Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics
Poster 3.6
Poster 3.6
Poster 3.6
Table 12: List of papers used in human evaluation. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.

### v3

@@ 5 Experimental Results and Analysis
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for human developers), as demonstrated by prior findings from Starace et al. [38]. Also, our goal is to provide a strong and faithful starting point that meaningfully aids reproduction efforts (supported by our results), rather than aiming for perfect reproduction out of the box. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution errors occur, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API calls or correcting data type mismatches, for successful execution (See Examples in Figures 6 and 7, alongside the detailed statistics in Table 9), which highlights that our generated repositories are near-executable with minimal human intervention.
6 Conclusion
In this work, we introduced PaperCoder, a framework that automatically generates code repositories from research papers in machine learning through a structured, three-stage pipeline. Specifically, we defined a high-level roadmap, system architecture, execution logic, and configuration via the planning stage, which are then enhanced through detailed per-file analysis, followed by the sequential code generation informed by artifacts from prior stages. To validate PaperCoder, we performed evaluations on two benchmarks: our Paper2CodeBench, comprising recent papers from top-tier machine learning venues, and (recently released) PaperBench Code-Dev, providing fine-grained evaluation protocols, on which PaperCoder consistently outperforms existing baselines on both model-based and human evaluations. Furthermore, additional analyses demonstrate its robustness and practicality: it remains effective across different LLM backbones, shows strong executability with only 0.81% of the lines requiring minor fixes, and benefits from each stage in the pipeline. We envision PaperCoder as one important step toward accelerating scientific progress by aiding the reproduction of research papers.
@@ A.1 Implementation Details
All experiments are conducted using the o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we use openreview_scraper [^4] https://github.com/pranftw/openreview_scraper with the OpenReview API [^5] https://docs.openreview.net/reference/api-v2 and the Semantic Scholar API [^6] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library [23] [^7] https://github.com/allenai/s2orc-doc2json. Notably, with o3-mini-high to generate repositories for 90 papers, the total API cost of PaperCoder amounts to \$76.65, resulting in an average cost of approximately \$0.90 per paper.
@@ A.2 Human Evaluation Process
Given the complexity of the task (requiring comprehension of scientific papers and their associated implementations), we recruit participants who have at least one peer-reviewed paper and a degree in computer science. We note that they were compensated at a rate of \$15 per hour. For annotation, they were provided with a 4-page document, which includes task instructions, annotation examples, and 10 generated repositories grouped into three sets, as follows: (Group 1) Model Variants of Our Method that includes repositories generated by our system using different backbone models (e.g., o3-mini vs. three open-source alternatives); (Group 2) Naive Baselines that includes repositories generated using only the Paper or the Abstract as input; and (Group 3) Related Works that includes repositories generated by existing software development frameworks, such as MetaGPT and ChatDev. Each repository was anonymized using a repo X naming format to prevent bias regarding the generation method. Following the question guidelines in the document, annotators reviewed and evaluated the repositories generated by different methods and models. Also, on average, evaluating 10 repositories for a single paper took approximately 45 minutes. Table 23 shows a detailed annotation example.
@@ A.3 Reference-Based Evaluation
In the reference-based evaluation setup, the repository may exceed the context length of (even frontier) LLMs. Following Starace et al. [38], when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
@@ A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the PaperBench Code-Dev benchmark used for evaluation requires an additional script file called reproduce.sh. To meet this requirement, we further prompt the coding agent to generate it for this benchmark and evaluate the code with it.
Appendix B Additional Experimental Results and Analysis
B.1 Code Availability
To estimate the proportion of accepted papers that release official code repositories, we collect data from three major machine learning conferences in 2024: ICLR, ICML, and NeurIPS. Specifically, we first retrieve the list of accepted papers from each conference using the OpenReview API [^8] https://docs.openreview.net/reference/api-v2 via openreview_scraper [^9] https://github.com/pranftw/openreview_scraper. While OpenReview abstracts sometimes include repository links, they are more commonly found in ArXiv [^10] https://arxiv.org/ abstracts. Therefore, we additionally use the Semantic Scholar API [^11] https://www.semanticscholar.org/product/api to obtain ArXiv abstracts corresponding to the accepted papers. We then check whether the abstract includes a GitHub URL as an indicator of released code. Table 7 summarizes the number of accepted papers, the number with publicly available repositories, and the corresponding percentages for each conference. On average, only 19.5% of accepted papers in them provide official code.
Table 7: Code availability across major machine learning conferences. We report the total number of accepted papers, the number of papers with publicly available code (identified via GitHub URLs in ArXiv abstracts), and the corresponding percentage for each venue. The last row shows the average across all three conferences.
Conference # of Accepted w/ Code Percentage (%)
ICLR 2024 2207 467 21.2
ICML 2024 2610 435 16.7
NeurIPS 2024 4006 825 20.6
Average 2941 576 19.5
Table 8: Average Replication Scores (%) on PaperBench Code-Dev. For all OpenAI models, the reasoning effort is set to high, and we take results for BasicAgent and IterativeAgent from the PaperBench paper [38]. For PaperCoder, we report the average and standard deviation over three runs, except for o1 and o3 due to costs.
Model Replication Score (%) Cost per Paper (\$)
BasicAgent (o3-mini) $5.1\pm 0.8$ N/A
BasicAgent (o1) $19.5\pm 1.2$ N/A
BasicAgent (claude-3-5-sonnet) $35.4\pm 0.8$ N/A
IterativeAgent (o3-mini) $16.4\pm 1.4$ N/A
IterativeAgent (o1) $43.3\pm 1.1$ 400.00
IterativeAgent (claude-3-5-sonnet) $27.5\pm 1.6$ N/A
PaperCoder (o3-mini) $45.14\pm 0.3$ 0.69
PaperCoder (o1) $38.31$ 8.81
PaperCoder (o3) $60.86$ 8.99
PaperCoder (claude-3-5-sonnet) $51.14\pm 1.4$ 3.61
B.2 PaperBench Code-Dev Results
We conduct additional experiments using various reasoning models, as shown in Table 8. Overall, our method achieves strong replication scores across models. Notably, when using o3, PaperCoder records the highest score of 60.86%. These results suggest that the latest and larger models, particularly those with stronger reasoning and coding capabilities, tend to yield better performance.
Appendix C Limitations and Future Work
While PaperCoder demonstrates strong performance in reproducing machine learning papers (where code implementations are particularly helpful and usually necessary for validating research ideas), its current scope is limited to this domain. Beyond this, we believe accelerating the reproduction of scientific discovery to other domains where code is not the primary medium for validation, such as theoretical mathematics, is an exciting direction for future work. In addition, the current version of PaperCoder processes only textual inputs, and extending it to process visual inputs (such as figures in papers) is an interesting avenue. Lastly, as with other repository-level code generation approaches, improving executability remains an important (but still challenging) direction for future work.
Appendix D Broader Impacts and Ethics Statement
Our work aims to generate faithful code repositories from scientific papers in machine learning, and we believe it has a substantial positive impact in contributing to open science and facilitating rapid experimentation. However, we also acknowledge potential risks and misuse of our framework. For example, some papers intentionally refrain from releasing implementations due to security concerns, such as those involving jailbreaking or exploitation techniques. Yet, our method could potentially be used to reproduce such sensitive implementations. To address such risks, in real-world production, it would be necessary to develop and incorporate safeguards (such as harmful content filters, protective prompting, and secure execution environments) to ensure responsible and safe use of our framework.
@@ Appendix D Broader Impacts and Ethics Statement
Modified lines (*.py) 2 0 10 26 10 8
Modified lines (config.yaml) 3 6 7 1 4 3.5
@@
Percentage 0.44 0.29 1.06 2.02 1.02 0.81
Table 10: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our framework as their top choice into six (described in the first row).
@@
16 13 8 6 7 4
Original 1# data_loader.py 2import openai 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9 10...(omited)... 11 12 response = openai.Completion.create( 13 engine=self.engine, 14 prompt=prompt, 15 max_tokens=150, 16 temperature=0.7, 17 n=1, 18 stop=["Task 16"] 19 ) 20 raw_text = response.choices[0].text.strip()
Editied 1# data_loader.py 2from openai import OpenAI 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9openai_api_key = os.environ.get("OPENAI_API_KEY") 10client = OpenAI(api_key=openai_api_key) 11 12...(omited)... 13 14 response = client.chat.completions.create( 15 model=self.engine, 16 messages=[{"role": "user", "content": prompt}], 17 max_completion_tokens=150, 18 temperature=0.7, 19 n=1, 20 stop=["Task 16"] 21 ) 22 raw_text = response.choices[0].message.content.strip()
Figure 6: Error case from Self-Instruct. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the version manually edited by the authors to correct the error. In this example, an outdated API call is updated to its latest version. In the original version, lines 2, 12, 14, 15, and 29 were removed; in the edited version, lines 2, 9, 10, 14, 16, 17, and 22 were added.
Original 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=lr) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model)
Edited 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=float(lr)) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model, trust_remote_code=True)
Figure 7: Error case from CoLoR. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the manually edited version by the authors. In this example, a numeric value is cast correctly and a required argument is added to enable execution. Lines 2 and 5 are modified in the corrected version.
Appendix E Prompts
@@
Prompt for analysis [System] You are an expert researcher, strategic analyzer and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to conduct a comprehensive logic analysis to accurately reproduce the experiments and methodologies described in the research paper. This analysis must align precisely with the paper’s methodology, experimental setup, and evaluation criteria. 1. Align with the Paper: Your analysis must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present your analysis in a logical, well-organized, and actionable format that is easy to follow and implement. 3. Prioritize Efficiency: Optimize the analysis for clarity and practical implementation while ensuring fidelity to the original experiments. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. REFER TO CONFIGURATION: Always reference settings from the config.yaml file. Do not invent or assume any values—only use configurations explicitly provided. [User]
@@
Figure 12: Prompts for analysis. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design.
@@
Appendix F Examples output of the planning phase
Figure 16: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (1).
Figure 17: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (2).
Figure 18: Artifact from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 19: Class diagram visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 20: Program call flow visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 21: Artifact from Step 1.3 (Logic Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
## config.yaml
training:
optimizer: "Adam"
beta1: 0.9
beta2: 0.98
epsilon: 1e-9
warmup_steps: 4000
total_steps: 100000 # Base model: 100K training steps (~12 hours on 8 P100 GPUs)
learning_rate_schedule: "d_model^-0.5 * min(step^-0.5, step * warmup_steps^-1.5)"
batch_tokens:
source: 25000
target: 25000
dropout: 0.1
label_smoothing: 0.1
model:
num_layers: 6
d_model: 512
d_ff: 2048
num_heads: 8
d_k: 64
d_v: 64
shared_embedding: true
data:
dataset: "WMT 2014 English-to-German"
encoding: "byte-pair encoding"
vocabulary_size: 37000
evaluation:
beam_size: 4
length_penalty: 0.6
max_output_length: "input_length + 50"
Figure 22: Artifact from Step 1.4 (Configuration File) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Table 11: List of ICLR 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
@@ A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Table 12: List of ICML 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
Oral 4
Autoformalizing Euclidean Geometry
Poster 4
Recurrent Distance Filtering for Graph Representation Learning
Poster 4
Poster 3.9
Token-level Direct Preference Optimization
Poster 3.9
Oral 3.8
CurBench: Curriculum Learning Benchmark
Poster 3.8
Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment
Poster 3.7
Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments
Poster 3.7
Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge
Poster 3.7
Poster 3.7
Poster 3.7
Poster 3.7
Poster 3.7
Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model
Oral 3.6
Oral 3.6
Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting
Oral 3.6
Poster 3.6
Poster 3.6
Graph Distillation with Eigenbasis Matching
Poster 3.6
Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?
Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics
Poster 3.6
Poster 3.6
Poster 3.6
Table 13: List of NeurIPS 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
@@
Table 14: List of papers used in human evaluation. We evaluate the official repository of each paper, released by the authors, using the model-based reference-free setting with gpt-4o-2024-11-20 as the evaluation model.
RepoName
@@
N/A
Silent Branding Attack: Trigger-free Data Poisoning Attack on Text-to-Image Diffusion Models
N/A
VideoRAG
VideoRAG: Retrieval-Augmented Generation over Video Corpus
3.0
RADA
3.0
STELLA_code
STELLA: Continual Audio-Video Pre-training with Spatio-Temporal Localized Alignment
3.3
prometheus-vision
Prometheus-vision: Vision-language model as a judge for fine-grained evaluation
3.1
CoLoR
Efficient Long Context Language Model Retrieval with Compression
3.0
Volcano
Volcano: Mitigating Multimodal Hallucination through Self-Feedback Guided Revision
3.2
N/A
T1: Tool-integrated Self-verification for Test-time Compute Scaling in Small Language Models
N/A